A spring has an unstretched length of 11 cm . It exerts a restoring force F when stretched to a length of 12 cm. For what total stretched length of the spring is its restoring force 3F? At what compressed length is the restoring force 2F?
Answer #1

F = k * delta(x)

F = k * (12 - 11) = k

k = F  

delta(x) = 1 cm ,

stretched length 11+1 = 12 cm

3F = (F) * delta(x)

delta(x) = 3 cm , stretched length 11+3 = 14 cm

2F = (F) * delta(x)

delta(x) = 2 cm , compressed length 11 - 2 = 9 cm